# Hybrid optimal control with mixed-integer Lagrangian methods

### v2

@@ 1 Introduction
University of the Bundeswehr Munich, Department of Aerospace Engineering, Institute of Applied Mathematics and Scientific Computing, 85577 Neubiberg, Germany. This research is funded by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr [EMERGENCY-VRD, MissionLab, MORE, SeRANIS]. dtec.bw is funded by the European Union – NextGenerationEU. Corresponding author: Alberto De Marchi.
We consider hybrid optimal control problems (OCP), that is, optimization problems involving possibly nonsmooth dynamics, mixed-integer states or controls, or logical constraints. Hybrid OCPs were investigated, e.g., by [15, 12] where hybrid dynamics are modeled with disjunctive programming. An important subclass of hybrid OCPs are mixed-integer OCPs with discrete-valued controls. Numerical methods often use variable time transformations, relaxation, or decomposition techniques; compare [10, 18, 2] for important developments. Such approaches are often tailored to discrete-valued controls and exploit the time structure.
@@
We study a tailored augmented Lagrangian (AL) method for the numerical solution of a class of MINLPs of the following type:
$\operatorname*{minimize}\quad$ $f(x)$ (P)
$\operatorname{subject~{}to}\quad$ $x\in\mathcal{X},\quad c(x)\in\mathcal{C}$
with $f\colon\mathcal{X}\to\mathbb{R}$ and $c\colon\mathcal{X}\to\mathbb{R}^{m}$ smooth functions, $\mathcal{C}\subset\mathbb{R}^{m}$ a nonempty closed convex set (projection-friendly in practice), and $\mathcal{X}$ a nonempty closed set with a mixed-integer linear (MIL) structure. In particular, we assume $\mathcal{X}$ to be described by linear inequalities and integrality constraints on some variables, that is,
$\mathcal{X}:=\left\{x\in\mathbb{R}^{n}\,\middle|\,\begin{array}[]{l}Ax=b,~{}x_{\textup{l}}\leq x\leq x_{\textup{u}},\\ x_{i}\in\mathbb{Z}\quad\forall i\in\mathcal{I}\end{array}\right\}$ (1)
for some matrix $A$, vectors $b$, $x_{\textup{l}}$ and $x_{\textup{u}}$, and an index set $\mathcal{I}\subseteq\{1,\ldots,n\}$. The general template (P) covers hybrid OCPs discretized with arbitrary schemes, and can capture various features such as mixed-integer control inputs, nonsmooth dynamics and state jumps.
@@
The approach here is to adopt the AL framework to solve the constrained problem (P) through a sequence of simpler subproblems, that is, moving the nonlinear constraints onto the objective. This methodology is motivated by sequential (partially) unconstrained minimization schemes [9], including (shifted) penalty and barrier (or interior point) methods, and by the availability of an affordable solver for problems of the form (P) without nonlinear constraints [6].
The work contributes a solution technique for (discretized) hybrid OCPs based on an affordable safeguarded AL approach with inexact subproblems. A summary of the optimality concept is described in Section 2, while algorithmic details are provided in Section 3. Hybrid OCPs are addressed with our method in Section 4, where numerical results demonstrate its modelling versatility and efficacy.
@@ 2.1 Mixed-integer polyhedral constraints
Recalling [6, Def. 2.3, Prop. 2.4], a first-order criticality measure associated to (2) is defined for all $\bar{x}\in\mathcal{X}$ and $\Delta\geq 0$ by
@@
As discussed in [6, §2.2], criticality is a necessary optimality condition for (2) and, as such, plays a crucial role when it comes to designing an affordable numerical method for solving (2). The denomination criticality, in contrast to stationarity, emphasizes that the former notion is in general stronger than the latter.
@@ 2.2 Nonlinear constraints
$\mathcal{L}(x,y):=f(x)+\langle y,c(x)\rangle.$ (4)
From the viewpoint of nonlinear programming, where local minima correspond to saddle points of the Lagrangian functions, we consider the following notion for KKT-like points of (P), building upon Definition 2.2.
@@ 3 Algorithmic Framework
cf. [5, §4]. Feasibility of (5) follows from $\mathcal{X}$ being nonempty whereas well-posedness is due to continuity of $\mathcal{L}_{\mu}(\cdot,\widehat{y})$ and can be guaranteed, e.g., by coercivity or (level) boundedness arguments.
Notice that $\mathcal{L}$ and $\mathcal{L}_{\mu}$ are smooth with respect to both primal and dual variables, thanks to smoothness of $f$, $c$ and convexity of $\mathcal{C}$, with derivatives
@@ 8       $\varepsilon_{j+1}\leftarrow\max\{\epsilon_{\textup{d}},\kappa_{\varepsilon}\varepsilon_{j}\}$
The major computational toll is taken by Algorithm 3.1, which aims to minimize $\mathcal{L}_{\mu_{j}}(\cdot,\widehat{y}^{j})$ over $\mathcal{X}$. In fact, only approximate criticality is required, and the sequential mixed-integer linearization scheme of [6, Algorithm 3.1] can be readily adopted to compute a suitable $\varepsilon_{j}$-critical point $x^{j}$, in the sense of Definition 2.2. Although the performance of the algorithm might be improved by warm-starting the inner solver and guaranteeing descent behaviour, these properties are not strictly required in the convergence analysis [4]. The classical first-order dual estimate update takes place at Algorithms 3.1 and 3.1. Given a (possibly inexact, first-order) solution $x$ to (5), the dual update rule in Algorithm 3.1 is designed toward the identity
@@
In analogy to magical steps [1, §8.2], [5, §4.1] and feasibility pumps [3], additional steps can be optionally included in Algorithm 3.1 to ease or accelerate the solution process. Such routines could pursue feasibility (switching to the minimization of an infeasibility measure) or refine the incumbent solution by solving a nonlinear program with fixed integer-valued variables [6, Rem. 3.1], [13, §5].
@@ 4 Numerical Results
ALM is initialized with zero dual estimate $y^{0}$, penalty parameter $\mu_{1}=10^{-2}$ and inner tolerance $\varepsilon_{1}=10^{-2}$. The choice of $x^{0}$ is based on a more sophisticated approach. Admissible points for the set $\mathcal{X}$ are computed as projections onto it, in the $\ell_{1}$-norm sense. Another routine was implemented to generate reasonable starting points and to refine solutions provided by ALM. While integer variables are relaxed in the former case, they are fixed (to an admissible value) in the latter case so that only real-valued variables are further optimized. In both scenarios, one ends up with a nonlinear problem, which is then tackled by Ipopt [19]. We use the numerical routine of [6] for executing Algorithm 3.1, with Gurobi for solving the MIL subproblems there.
@@ 4.1 Car with hysteretic turbo charger and drag
The first problem handled by the proposed method is inspired by [14, §V.A] and [6, §4.1]. The aim is to perform point-to-point one-dimensional motion planning of a car with hybrid nonlinear dynamics, extending previous models to accommodate a drag force term. The underlying double-integrator point-mass model is equipped with a hysteretic turbo accelerator. The car is described by its time-dependent position $x(t)$, velocity $v(t)$ and a turbo state $w(t)\in\{0,1\}$ for each time point $t\in[0,T]$, $T>0$. The control variables correspond to the inputs to the acceleration $a(t)$ and brake $b(t)$ pedals. The turbo mode becomes active or inactive once the velocity exceeds $v^{+}:=10$ or falls below the threshold $v^{-}:=5$, respectively. An activated turbo mode entails an effective increase of the nominal thrust by three times but leaves the braking force unaffected. These dynamics read
@@ 4.2 Lotka-Volterra fishing with total variation
for tracking the reference value $x_{r}=(1,1)$, starting from the initial state $x_{1}(0)=0.5$, $x_{2}(0)=0.7$ and with final time $T=12$ [18, §5].

### v3

@@ 1 Introduction
The authors are with the University of the Bundeswehr Munich, Department of Aerospace Engineering, Institute of Applied Mathematics and Scientific Computing, 85577 Neubiberg, Germany.This research was partially funded by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr [EMERGENCY-VRD, MissionLab, MORE, SeRANIS]. dtec.bw is funded by the European Union – NextGenerationEU.This work has been accepted for presentation at the 13th IFAC Symposium on Nonlinear Control Systems (NOLCOS 2025) and for publication in the IFAC-PapersOnLine proceedings.
We consider hybrid optimal control problems (OCPs), that is, optimization problems involving possibly nonsmooth dynamics, mixed-integer states or controls, or logical constraints. Hybrid OCPs were investigated, e.g., by [13, 11] where hybrid dynamics are modeled with disjunctive programming. An important subclass of hybrid OCPs are mixed-integer OCPs with discrete-valued controls. Numerical methods often use variable time transformations, relaxation, or decomposition techniques; compare [9, 17, 2, 14] for important developments. Such approaches are often tailored to discrete-valued controls and exploit the time structure.
@@
We study a method for the numerical solution of a class of MINLPs of the following type:
$\operatorname*{minimize}~{}$ $f(x)$ $\operatorname{subject~{}to}~{}$ $x\in\mathcal{X},\quad c(x)\in\mathcal{C}$ (P)
with $f\colon\mathcal{X}\to\mathbb{R}$ and $c\colon\mathcal{X}\to\mathbb{R}^{m}$ differentiable functions, $\mathcal{C}\subset\mathbb{R}^{m}$ a nonempty closed convex set (projection-friendly in practice), and $\mathcal{X}$ a nonempty closed set with a mixed-integer linear (MIL) structure. In particular, we assume $\mathcal{X}$ to be described by linear inequalities and integrality constraints on some variables, that is,
$\mathcal{X}:=\left\{x\in\mathbb{R}^{n}\,\middle|\,\begin{array}[]{l}Ax\leq b,\\ x_{i}\in\mathbb{Z}\quad\forall i\in\mathcal{I}\subseteq\{1,\ldots,n\}\end{array}\right\}$ (1)
for some matrix $A$, vector $b$ and index set $\mathcal{I}$. The general template (P) covers hybrid OCPs discretized with arbitrary schemes, and can capture various features such as mixed-integer control inputs, nonsmooth dynamics and state jumps.
@@
This work proposes a solution technique for (discretized) hybrid OCPs based on a safeguarded augmented Lagrangian (AL) approach, traditionally confined to nonlinear programming. Adopting the AL framework of [1], the constrained problem (P) is solved through a sequence of simpler subproblems. This methodology is motivated by sequential (partially) unconstrained minimization schemes, including penalty and barrier methods; see [8]. The AL subproblems are tackled by the tailored solver proposed in [6]. Such a practical AL algorithm for MINLPs is our main contribution, intended to seek good local, possibly non-global, solutions to (P).
The remainder of this work is structured as follows. A summary of the optimality concept is described in Section 2, while algorithmic details are provided in Section 3. Hybrid OCPs are addressed with our method in Section 4, where numerical results demonstrate its modelling versatility and efficacy.
@@ 2.1 Mixed-integer polyhedral constraints
Recalling [6, Def. 2.3, Prop. 2.4], a first-order optimality measure associated to (2) is defined for all $\bar{x}\in\mathcal{X}$ and $\Delta\geq 0$ by
@@
As discussed in [6, §2.2], criticality is a necessary optimality condition for (2) and, as such, plays a crucial role when it comes to designing an affordable numerical method for solving (2). The denomination criticality, in contrast to stationarity, emphasizes that in general the former notion implies (hence, is stronger than) the latter.
@@ 2.2 Nonlinear constraints
$\mathcal{L}(x,y):=f(x)+\langle y,c(x)\rangle$ (4)
where $y$ denotes a Lagrange multiplier. From the viewpoint of nonlinear programming, where local minima correspond to saddle points of the Lagrangian functions, we consider the following notion for KKT-like points of (P), building upon Definition 2.2.
@@ 3 Algorithmic Framework
cf. [5, §4]. Feasibility of (5) follows from $\mathcal{X}$ being nonempty whereas well-posedness is due to continuity of $\mathcal{L}_{\mu}(\cdot,\widehat{y})$ and can be guaranteed, e.g., by coercivity or (level) boundedness arguments. Notice that $\mathcal{L}$ and $\mathcal{L}_{\mu}$ are differentiable with respect to both primal and dual variables, thanks to convexity of $\mathcal{C}$, with derivatives
@@
Furthermore, note that the input requirement in Algorithm 3.1 is not a hard restriction: given some $x^{0}\notin\mathcal{X}$, it is enough to project it onto $\mathcal{X}$ by solving a MILP, for instance $\min_{x\in\mathcal{X}}\|x-x^{0}\|_{1}$.
@@ 8       $\varepsilon_{j+1}\leftarrow\max\{\epsilon_{\textup{d}},\kappa_{\varepsilon}\varepsilon_{j}\}$
The major computational toll is taken by Algorithm 3.1, which aims to minimize $\mathcal{L}_{\mu_{j}}(\cdot,\widehat{y}^{j})$ over $\mathcal{X}$. In fact, only approximate criticality is required, and the sequential mixed-integer linearization scheme of [6, Alg. 3.1] can be readily adopted to compute a suitable $\varepsilon_{j}$-critical point $x^{j}$, in the sense of Definition 2.2. Although the performance of the algorithm might be improved by warm-starting the inner solver and guaranteeing descent behaviour, these properties are not strictly required in the convergence analysis [4]. The classical first-order dual estimate update takes place at Algorithms 3.1 and 3.1. Given a (possibly inexact, first-order) solution $x$ to (5), the dual update rule in Algorithm 3.1 is designed toward the identity
@@
In analogy to magical steps [1, §8.2], [5, §4.1] and feasibility pumps [3], additional steps can be optionally included in Algorithm 3.1 to ease or accelerate the solution process. Such routines could pursue feasibility (switching to the minimization of an infeasibility measure) or refine the incumbent solution by solving a nonlinear program with fixed integer-valued variables [6, Rem. 3.1], [12, §5].
@@ 4 Numerical Results
ALM is initialized with zero dual estimate $y^{0}$, penalty parameter $\mu_{1}=10^{-2}$ and inner tolerance $\varepsilon_{1}=10^{-2}$. The choice of $x^{0}$ is based on a more sophisticated approach. Admissible points for the set $\mathcal{X}$ are computed as projections onto it, in the $\ell_{1}$-norm sense. Another routine was implemented to generate reasonable starting points and to refine solutions provided by ALM. While integer variables are relaxed in the former case, they are fixed (to an admissible value) in the latter case so that only real-valued variables are further optimized. In both scenarios, one ends up with a nonlinear problem, which is then tackled by Ipopt [18]. We use the numerical routine of [6, §4] for executing Algorithm 3.1, with Gurobi for solving the MIL subproblems there.
@@ 4.1 Car with hysteretic turbo charger and drag
The first problem handled by the proposed method is inspired by one from [6, §4.1]. The aim is to perform point-to-point one-dimensional motion planning of a car with hybrid nonlinear dynamics, extending the previous model to accommodate a drag force term. The underlying double-integrator point-mass model is equipped with a hysteretic turbo accelerator. The car is described by its time-dependent position $x(t)$, velocity $v(t)$ and a turbo state $w(t)\in\{0,1\}$ for each time point $t\in[0,T]$, $T>0$. The control variables correspond to the inputs to the acceleration $a(t)$ and brake $b(t)$ pedals. The turbo mode becomes active or inactive once the velocity exceeds $v^{+}:=10$ or falls below the threshold $v^{-}:=5$, respectively. An activated turbo mode entails an effective increase of the nominal thrust by three times but leaves the braking force unaffected. These dynamics read
@@ 4.2 Lotka-Volterra fishing with total variation
for tracking the reference value $x_{r}=(1,1)$, starting from the initial state $x_{1}(0)=0.5$, $x_{2}(0)=0.7$ and with final time $T=12$ [17, §5].